### v2

# Two Sides of the Same Coin: Large-scale Measurements of Builder and Rollup after EIP-4844

Abstract.
Web3 is reshaping decentralized ecosystems through innovations like Ethereum. Recently, EIP-4844 is implemented in Ethereum to support its Layer-2 scaling solutions, which introduces a new 128 KB data structure called blob. This upgrade incorporates type-3 transactions with blobs to verify data availability and reduce gas costs for rollups, significantly affecting the strategies of both builders and rollups. In this paper, we present an in-depth study of emerging strategies in builder and rollup markets after EIP-4844, containing hundred million transactions. We find that the efficiency of builder and rollup strategies is interdependent, akin to two sides of the same coin—both cannot be optimized simultaneously. That is, when builders operate efficiently, rollups tend to overpay in fees, conversely, when rollups optimize their costs, builders may incur losses in inefficient transaction selection. From the side of builders, our results show that 29.48% of these blocks have been constructed inefficiently, which does not produce sufficient profits for builders. Through our evaluation from the side of rollups, we find that over 72.53% of type-3 transactions pay unnecessary fees, leading to notable economic costs of rollups. Our work provides critical insights into optimizing block construction and transaction strategies, advancing the economic efficiency and data scalability of Web3 infrastructures, yet, much like balancing a seesaw, the efficiency of builders and rollups cannot be optimized concurrently.
Ethereum, Layer-2, Builder, Rollup, EIP-4844, Blob
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the Conference 2025; June 28–October 02, 2018; NY, USAisbn: 978-x-4503-XXXX-X/18/06ccs: General and reference Empirical studiesccs: General and reference Measurementccs: General and reference Evaluation
1. Introduction
Web3 leverages decentralized technologies such as blockchain to enable peer-to-peer interactions without intermediaries. A groundbreaking and transformative application of decentralized technology is Ethereum, an open and decentralized blockchain platform that facilitates the development and execution of smart contracts and decentralized applications (Gilbert, 2022; Wang et al., 2022b).
In the Ethereum mainnet, the Proposer Builder Separation (PBS) mechanism was proposed to separate the role of the miner by introducing the builder and the proposer (Buterin, 2021a). The task of builders is to construct the block, where they sort the transactions to maximize the extractable value and submit the block to proposers (Daian et al., 2020; Heimbach and Wattenhofer, 2022). Each transaction within the block originates from two sources: the public mempool, which is accessible to everyone, and private channels that deliver transactions directly to specific builders (Qin et al., 2022).
Ethereum Layer-2 refers to scaling solutions built on top of the Ethereum mainnet to enhance its performance and throughput, with rollups being one of the most prominent approaches that bundle multiple transactions into a single batch before submitting them to the mainnet. Ethereum has implemented EIP-4844 in the Dencun upgrade to improve network scalability and lower transaction costs (Ethereum.org, 2024a). This upgrade presents blob, a new data format of 128 KB designed to facilitate rollups on Ethereum’s Layer-2, alongside blob-carrying transactions known as type-3 transactions (Vitalik Buterin, 2023). By offloading data storage to blobs, the gas fee is significantly reduced for rollups. Type-3 transactions are much larger and therefore take more space in blocks (Park et al., 2024; Gomez, 2024b). Blocks with large size take longer to propagate through the network. On the one side, builders need to balance the block size with the profit they can obtain (Wahrstätter, 2024b; Gomez, 2024a). They also need to decide whether to choose type-3 transactions in their blocks. This results in different behaviors among builders in response to type-3 transactions. On the other side, rollups need to issue their blobs into type-3 transactions and wait for being included in blocks. Type-3 transactions always carry the different number of blobs and different fees, which indicate various strategies of rollups (Cui, 2024). Evidently, the strategies among rollups are also complicated.
We evaluate the strategies of builders and rollups after EIP-4844. We find that the efficiency of their strategies is like the two sides of the same coin. When the builders strategy is efficient, rollups may need to pay more fees. Oppositely, when rollups adopt the efficient strategy and offer lower fees, builders may face block profits lower than potential profits. In the side of builders, we first examine the new strategy that builders employed after EIP-4844 (Section 4). Although the average total block size has been increasing to nearly 400 KB, the size of transactions from public mempool in the block has decreased from 150 KB before the upgrade to about 30 KB. Builders have taken to shrinking the size of the rest of the block to contain type-3 transactions, which prompts us to delve deeper into the efficiency of this strategy. We recognize that the builder strategy is not efficient enough after EIP-4844. In the side of rollups, they need to pay higher fees to ensure inclusion in efficient blocks. However, we find that most type-3 transactions issued by rollups pay too high fees (Section 5). Our primary contributions are:
  1. (1)
    We build a large-scale dataset to quantify builders and rollups market in Ethereum from after EIP-4844 to August, which includes 319,529,950 transactions with 1,336,822 type-3 transactions. To the best of our knowledge, we are the first to illuminate different strategies and efficiencies of builders and rollups in the network after EIP-4844. Our dataset provides the most complete labelling of builders and rollups so far.
  2. (2)
    We identify the strategies that different builders adopt after EIP-4844. By examining the size of transactions versus the profit paid to builders, we provide a model for measuring efficient blocks for builders. We find that 29.48% of the blocks containing type-3 transactions do not give the builder enough profits to compensate for the loss of excluding other transactions. This will assist builders in improving their strategies for constructing blocks for higher profit margin.
  3. (3)
    We explore the sending patterns of different rollups for blob and type-3 transactions and find that the behavior within the market is distinctive. We measure efficient type-3 transactions for rollups and observe that 72.53% of type-3 transactions in the market are inefficient. This work contributes to how rollups can optimize their pricing of type-3 transactions to achieve cost savings.
  4. (4)
    Furthermore, we evaluate the losses of other flawed strategies in the market and reveal significant losses of 186.92 ETH for rollups. We also characterize the impact of one misuse of blobs. It induces a delay of over 19 seconds in the type-3 transactions of rollups.
2. Backgrounds
[…]
Illustration of the block building process in Ethereum.
Figure 1. Illustration of the block building process in Ethereum. Builder is divided into MEV-Boost builders and Vanilla builders according to their transaction source. Their block would contain type-3 transactions introduced by EIP-4844.
In this section, we provide an overview of Ethereum builders, rollups and EIP-4844.
2.1. Ethereum
Ethereum stands as a leading decentralized blockchain platform (Buterin et al., 2013). With an upgrade known as the Merge (Foundation., 2023), Ethereum has transitioned from a PoW consensus, which is energy intensive, to a PoS mechanism. The Proposer Builder Separation (PBS) mechanism has been proposed as an important development after the Merge (Ethereum.org, 2023). The traditional role of validators, formerly known as miners, is divided into separate entities: block builders and block proposers (i.e., the validators themselves) (Buterin, 2021a, b). Flashbots introduced MEV-Boost as an intermediate measure to enable the implementation of PBS (Wahrstätter, 2023). The other roles are searchers and relays. The role of searchers is to identify and leverage profit extraction opportunities in the public mempool (Qin et al., 2022; Wang et al., 2022a; Zhou et al., 2023). Once they finish extracting the value, they would send the transaction bundles to builders via private channels (Mamageishvili et al., 2024). Relays are introduced to establish the trust required between block builders and proposers (Flashbots, 2022).
As shown in Figure 1, builders receive two types of transactions, public transactions from public mempool (Wackerow, 2024a) and private transactions from private channels (Raun et al., 2023). In addition to the transaction priority fee and coinbase transafer, the builder may profit by collaborating privately with the transaction sender (EigenPhi, 2024). Builders prioritize transactions with higher priority fee and subsequently rank transactions by their potential revenue, before determining which transactions to include in blocks. This creates a block space race in which users can bid for faster inclusion through higher priority fees (Liu et al., 2022). Users who want faster confirmations can offer higher priority fees, while those willing to wait can pay just the base fee.
Currently, it is optional for Ethereum validators to participate in PBS (Wahrstätter, 2024b). Validators who opt not to use PBS and instead build blocks independently are known as Vanilla builders. Vanilla builders receive no private order flow because they are not specialized in block building tasks. They complete both block building and validation tasks simultaneously and currently account for 10% of all builders in the network (Wahrstätter, 2023).
2.2. Rollup
Rollups are Layer-2 scaling solutions designed to improve the efficiency and throughput of the Ethereum network by offloading transaction processing from the Ethereum mainnet. To protect the data availability of Ethereum (Pomerantz, 2024), rollups work by bundling or rolling multiple transactions into a single batch. After processing, the rollup smart contract on Ethereum updates the state with a single, compressed transaction that includes the aggregated data of all individual transactions. It will be submitted to the mainnet, thereby reducing the computational load and gas fee for each individual transaction. Rollups maintain the security and decentralization of Ethereum while providing faster and cheaper transactions, making them a crucial component of the Ethereum scalability roadmap (Smith, 2024).
Rollups come in two main types: optimistic rollups (Guimarães, 2024) and zero-knowledge rollups (Wackerow, 2024b). Optimistic rollups assume transactions are valid and only run fraud proofs when discrepancies are detected, whereas zero-knowledge rollups use cryptographic proofs to validate the correctness of the bundled transactions. Both solutions have unique trade-offs and use cases, but share the common goal of making Ethereum more scalable without sacrificing security.
2.3. EIP-4844
Rollups used calldata to store state roots and proofs (Pomerantz, 2024), ensuring that the rollup state could be re-established even if a transaction was challenged. Calldata is the payload contained in a transaction that can be used to store arbitrary information on the chain. While calldata is cost-effective for smaller amounts of data, it is permanently stored in Ethereum state and is therefore relatively expensive for larger data requirements. The high cost of calldata storage becomes a bottleneck for scaling.
In order to improve the scalability and data availability of the network, Ethereum proposed a new sharding design, Danksharding (Ethereum.org, 2024b). Danksharding is a unified sharding design that differs from traditional sharding by managing data storage and availability in a more streamlined way. Instead of increasing the space for transactions, Ethereum sharding provides more space for blobs of data, which the Ethereum protocol itself does not interpret. EIP-4844, also known as Proto-Danksharding, implements the same transaction formats and validation rules as the new framework and logic of Danksharding. EIP-4844 was launched in March as a core part of the Dencun upgrade in 2024 (Ethereum.org, 2024a), thereby facilitating the network’s preparation for full Danksharding by allowing the inclusion of large data in transactions (Pomerantz, 2024). The EIP-4844 represents a significant milestone in Ethereum’s roadmap (Ethereum.org, 2024c), aiming to improve the scalability, security, and usability of the network.
There are three types of transaction before EIP-4844, type-0, type-1 and type-2 transaction, and type-2 transactions always bring priority fee for inclusion by builders quickly. EIP-4844 focuses on type-3 transactions (Park et al., 2024) which are similar to regular transactions like type-2 transactions, but carry an additional data component called ‘blob’. These oversized blobs (128 KB) are much cheaper than the equivalent amount of calldata because they are not permanently stored on the Ethereum (Pomerantz, 2024), and rollups can also send their type-3 transactions from public mempool or private channels. This new type of transaction meets the growing need for efficient data processing. The blob data remains inaccessible during the execution of the Ethereum virtual machine, and the validator can only see the commitment to the blob, but these blobs can be used for Layer-2 scaling solutions and other off-chain protocols that require large amounts of data throughput (Gogol et al., 2024a).
3. Dataset
We generate our data in two stages to study the dynamic building and rollup market. First, to support the analysis, we accessed public on-chain data, including block, transaction, and receipt data. To collect transactions in the public mempool, we use archived data from Flashbots. We decode the data that is used to identify block builders to label both MEV-Boost builders and Vanilla builders, and collect rollup address to identify rollup applications and others. As far as we are aware, this is the first time such an extensive data set with labels has been constructed. To facilitate further research, our labeling dataset would be released in the published version.
3.1. Data Collection
On-chain Data.  To build the on-chain dataset, we deploy the Ethereum execution client Erigon (Ledgerwatch, 2024) and the consensus client Lighthouse (Prime, 2024). We collect on-chain data covering the period from December 1, 2023 to August 31, 2024. The on-chain dataset that we generated contains 1,967,415 blocks, 319,529,950 transactions and receipts. The dataset we built also includes the hash address of each blob in 1,336,822 type-3 transactions.
Public Mempool Data.  We next acquire the public mempool archive data from Flashbots Mempool Dumpster (Flashbots, 2023), which is updated every day. The mempool data provides the transaction inclusion delay and the information such as the type and size of a transaction (Hager, 2022). Note that the transaction inclusion delay is calculated from the time interval between the first time the transaction is detected and the time when it is included in a block. The public mempool data range from the same period of on-chain dataset and have 318,175,240 public transactions.
3.2. Labeling
Builder Labeling.  To better match the builder and its blocks, we decode the extra_data from the hexadecimal string to the origin value and manually check the label of each vanilla builders. We compare our builder labeling dataset with some existing public dataset (Network, 2023) created using builders’ public key, and our dataset can cover vanilla builders which was ignored in other datasets. Our labels contain 56 MEV-Boost builders and 318 Vanilla builders.
Rollup Labeling.  To identify different rollups in Layer-2, we use the address disclosed by (hildobby, 2024; Tran, 2024) to match the corresponding rollups. As some rollups would add their chain ID to the to address, to label more rollups in the network, we extract the chain ID of these rollups, requesting the chain ID towards the API from Chainlist. [^1] https://chainlist.org/ The address we gathered covers 35 main rollups. We also compare each address with official documents to ensure that our dataset is accurate. To our knowledge, our rollup labels are by far the most covered. Additionally, we also find a large amount of type-3 transactions are used for blobscriptions (Blobscriptions, 2024), to make it clear, we use the API from Ethscription. [^2] https://api.ethscriptions.com/v2/ethscriptions/ Utilizing this API, it is possible to retrieve a blobscription via its transaction hash. We identify and label all instances within type-3 transactions belonging to blobscriptions. Consequently, our labeling process encompasses 106,051 type-3 transactions used for blobscriptions. We identify all unlabeled addresses in our type-3 transaction dataset and annotate 5,655 Blobscription addresses following EIP-4844.
4. Characterizing Builder Behaviors
It is crucial to understand whether builders can efficiently select high-value transactions after EIP-4844. In this section, we first quantify the impacts of introducing large-volume type-3 transactions on builders’ construction strategies. We then provide a model to access the builders’ efficient strategy and analyze how the efficiency of the builder market changes.
4.1. Block Building Strategies
(a) Block size.
[…]
Block size.
(b) Size excluding type-3 transactions.
[…]
Type-3 excluded block size.
(c) Size from public mempool excluding type-3 transactions.
[…]
Block size from public mempool.
Figure 2. The average size in block among builders
[…]
Average block size from builders.
To demonstrate the difference of builder’s strategy towards block construction, we analyze the block size alongside changes in specific components. We present the average block size, the size of non type-3 transactions within the block, and the size of non type-3 transactions from the public mempool within the block in 2(c), respectively, covering the period from the beginning of December 2023 to the end of August 2024, offering some valuable insights. Firstly, in 2(a), the black line represents the overall average block size, which is generally on an upward trend. EIP-4844 leads builder to switch their strategy and the total size of the block increases significantly, from less than 150 KB to near 400 KB, which is roughly a 2.5-fold increase. Additionally, the blue and green lines represent the Flashbots builder and the rsync builder, respectively, where Flashbots changed its strategy in May, choosing to exclude type-3 transactions from the block, and then continuing to include them in July.
To measure the builder’s varying behavior towards other transactions after the upgrade, in 2(b) we observe the changes in the size of the non type-3 transaction portion of the block, which is the space occupied by all non type-3 transactions within a block. Clearly, after exceeding 300 KB in January, the average block size without type-3 transactions began to decline to 80 KB by June, reducing by almost four times. The pink and yellow lines represent the strategy of the Beaverbuild builder and Vanilla builders, who were the first to implement the strategy of decreasing the total size of other transactions within a block following EIP-4844. Nearly 60 percent of the market share is held by these builders, and the majority of the blocks they constructed are below the average size prior to May. By March, Vanilla builders transitioned to creating smaller blocks of other transactions, possibly due to not receiving a higher volume of orders and lacking the professional expertise to build blocks. Notably, starting from June, the top five market share builders, including Flashbots builder and Titan builder who previously built the largest blocks, have begun creating smaller sized blocks. In conjunction with 2(a), this finding suggests that the builders are reducing the size of the other parts in order to include the type-3 transaction in the block. It is worth noting that the volume size reduces to below 80 KB after June, and remains stable at around 80 KB thereafter. That is, the increase in type-3 transmissions has squeezed space for type-2 transactions.
The non type-3 transaction sizes within a block, as depicted in 2(b), can be categorized into two channels: public and private. Due to the exclusive access of the private channel to specific builders, in 2(c), we illustrate the size of other transactions from the public mempool within a block to further analyze the strategy behind the smaller size. The general trend observed in 2(c) is largely similar to that of 2(b). This suggests that the reduction in block size is due to builders minimizing the size of transactions taken from the public mempool. It is concerning to recognize that all builders employ a detrimental tactic to select small-volume transactions from the public mempool, potentially causing extended delays for numerous transactions that should have been included in the block. The detailed analysis of the result is provided in Appendix F.
4.2. Efficient Strategy for Builders
[…]
Figure 3. The profit of type-3 transactions and public mempool over time. The left y-axis shows the potential profit and right y-axis represents the profit difference.
For builders, the decision between type-2 and type-3 transactions constitutes a significant consideration. Both transaction types provide builders with profits via the ‘maxPriorityFeePerGas’ field. We show in the previous section that builders consciously reduce the volume of type-2 transactions and choose to select more type-3 transactions. To identify the efficient strategy of the choice, we derive the model of the block floor profit by using type-2 transactions in the public mempool. The efficient strategy for builders is to choose type-3 transactions when their profit is above the block floor profit.
We denote a set of type-2 transactions as $\mathbb{T}=\{tx_{1},tx_{2},\ldots,tx_{u}\}$ and a set of type-3 transactions as $\mathbb{B}=\{b_{1},b_{2},\ldots,b_{k}\}$ in the time frame of each block construction. The number $u$ is the quantity of type-2 transactions, and the number $k$ is the quantity of type-3 transactions in this time frame. The set of transactions in the public mempool is represented as $\mathbb{T}_{p}$, $\mathbb{T}_{p}\in\mathbb{T}$. The set of included transactions is represented as $S_{in},S_{in}\subseteq\mathbb{T}$. Therefore, type-2 transactions remaining in the public mempool are represented as $\mathbb{T}_{r}=\mathbb{T}_{p}\setminus S_{in}$.
For builders, the research shows that the factors they are concerned about are the profit of builders and the block size (Wu et al., 2024; Wahrstätter et al., 2023). Therefore, for each transaction $tx_{i}$ in $\mathbb{T}$, we denote its size as $s_{i}$ and its profit of builders as $v_{i}$. Then the transaction $tx_{i}$ can be further denoted as $tx(s_{i},v_{i})$. For each type-3 transaction $b_{j}$ in $\mathbb{B}$, we denote the number of blobs included in it as $q_{j}$ and its profit of builders as $p_{j}$. As specified by EIP-4844, each blob has a size of 128 KB. Therefore, the type-3 transaction $b_{j}$ can be further represented as $b(128q_{j},p_{j})$.
Block Floor Profit. The block floor profit represents the maximum profit that can be achieved by selecting from $\mathbb{T}_{r}$. The total size of these selected type-2 transactions is subject to the number of blobs. Formally, the block floor profit $F_{tx}(b)$ of type-3 transaction $b$ can be defined as
(1) $F_{tx}(b)$ $=\max\sum_{tx_{i}\in\mathbb{T}_{r}}v_{i},$
[…]
Intuitively, it represents the maximum profit that builders can obtain from the remaining type-2 transactions in the public mempool to compensate for the consumption of block size by type-3 transactions.
Efficient Block. The block floor profit represents the highest profit that builders can generate from type-2 transactions left in the public mempool, subject to the size of blobs. So for builders, the efficient strategy is to choose type-3 transactions if the profit satisfy
(2) $\sum_{b_{j}\in S_{in}}p_{j}\geq F_{tx}({b}).$
And if Equation 2 does not hold, then the efficient strategy is to directly select type-2 transactions from the public mempool. Blocks satisfying the efficient strategy are defined as efficient blocks. We find that there are still many inefficient blocks. Specific data and analysis will be provided in the next section.
4.3. The Efficiency of Builders
Table 1. The overall inefficient block distribution over time.
Month Inefficient Total Inefficient Proportion
March 26,450 47,604 55.56%
April 41,779 84,631 49.37%
May 47,351 105,382 44.93%
June 31,540 139,529 22.60%
July 27,965 145,726 19.19%
August 23,139 149,515 15.48%
Overall 198,224 672,388 29.48%
To measure whether builder new strategies are efficient or not, we employ a greedy algorithm considering size and profit to measure whether the priority fee given to each blob can match the priority fee that can be extracted in the public mempool during the concurrent period. Note that before being executed, only the maximum priority fee, i.e. maxPriorityFeePerGas field, can be seen. It represents the profit for builders in one transaction. We use this value to calculate the potential profit for builders. To filter out transactions that builders have simulated as likely to fail, we filter out transactions in the public mempool with a delay of more than 100 seconds to make the calculation more accurate. The pseudocode of the algorithm is provided in Appendix C.  Figure 3 plots the profit that builders actually extract from type-3 transactions versus the profit that can be extracted from public mempool in the same period from March to August. We compute the profit difference between the profit of type-3 transactions and the profit that can be extracted from the public mempool, and indicate this value with red and yellow bars. The red bar indicates that this difference is negative, which suggests that the builders could have made more profit from the public mempool, but they chose to include the less profitable type-3 transactions in the block. Conversely, if the colour is yellow, it indicates that the value is positive and the builder made more profit from type-3 transactions. It is clear that after EIP-4844, the builder misses the significant value of the public mempool. As Figure 3 shows, in March, builders can potentially make more than five times as much profit from the public mempool as they can from type-3 transactions. However, this loss diminishes over time and reverses in July, when the builder makes more profit from type-3 transactions than it is able to make from the public mempool.
[…]
Figure 4. The proportion of inefficient blocks per builder. The number above the bar is the total blocks specific builder built.
Table 1 illustrates the details from March to August, including the amount of efficient and inefficient blocks and their proportion Of the 636,981 blocks containing type-3 transactions. We find that 196,090 blocks are inefficient, which is approximately 29.48%. Fortunately, while the percentage of inefficient blocks was 55.56 percent after EIP-4844, by August it was only 15.48 percent, suggesting that the block building market is gradually becoming more efficient.  Figure 4 shows the ratio of efficient blocks among major builders during the same period. The y-axis shows the proportion of efficient and inefficient blocks. Beaverbuild has a market share of almost 50%, and is the most efficient compared to other builders. In contrast, non-specialist builders, vanilla builders who do not specialize in block construction, have nearly 30% more inefficient blocks than beaverbuild, and more than half of the blocks they construct are inefficient.
5. Understanding Rollup Strategies
The efficiency of the strategy between builder and rollup is like two sides of the same coin. Efficient builders typically result in rollups facing higher fees, whereas efficient rollups can impose losses on builders when selecting transactions. In this section, we examine the opposite side of the market, rollup’s strategy. We start by analyzing the different rollup behaviors after EIP-4844. We delineate the efficient strategy employed by rollups and conduct an analysis about the efficiency of type-3 transactions. We also analyze two flawed practices and measure their impact.
5.1. Type-3 Transactions Usage
[…]
Figure 5. The landscape of type-3 transactions after EIP-4844. The y-axis represents the daily usage and different colors indicates various rollups.
After the upgrade, in total 1,336,822 type-3 transactions have been sent to the Ethereum mainnet. Figure 5 visualizes the daily share of type-3 transactions by the dominant rollups and a specific role after EIP-4844, in terms of the total number of blob-carrying transactions. We notice that the number of transactions for each rollups varies over time, with some rollups showing significant spikes or drops. One of the most noticeable spikes is in Blobscriptions, which began hitting the market at the end of March, reached its peak, and then began to decrease dramatically. The purpose of Blobscriptions is not to validate rollup data, but rather to publish it. Other rollups show varying increases and decreases in usage, which is related to the choices made by each rollup application. Obviously, the usage of type-3 transactions in general has been increasing after EIP-4844, with more and more rollup applications starting to use blobs to validate data. Instead of immediately using blob to transfer data, some rollups (e.g., Taiko) only started adopting type-3 transactions in May of this year. However, there are still many rollup applications (such as Arbitrum and Base) that have been using type-3 transactions since EIP-4844.
[…]
Figure 6. The blob packing strategies among rollups.
In  Figure 6, we plot the blob packing strategy for each mainstream rollup from 13 March to 31 August. The different colours represent the number of blobs carried in type-3 transactions and the Y-axis represents the total number of transactions. It is clear that rollup makes different choices in selecting the number of blobs carried in a type-3 transaction. Of the 343,468 transactions issued by Taiko, it puts only one blob into the transaction, and the same strategy is used by Scroll, Starknet, and Zircuit.We further investigate the reasons for this in Appendix B, and find that the latency of the transactions that carry only one blob is the lowest. Therefore, it is likely that this strategy is optimising the latency of the rollup application so that the user’s transactions can be verified the fastest. In contrast, Base and Op Mainnet, which have been using blobs to send data, put as many blobs as possible into a transaction, and most of their transactions carry five to six blobs in a single transaction.In addition, there are some rollups that do not limit how many blobs they carry in a single transaction, such as Linea and Blast. We also show in Appendix A the percentage of type-3 transactions issued by different rollups that carry the same number of blobs. 69.85 % of the transactions in the network carry only one blob, while a staggering 1,239 transactions carry four blobs.
5.2. Efficient Strategy for Rollups
Table 2. The overall inefficient type-3 transactions distribution.
Month Inefficient Total Inefficient Proportion
March 52,298 106,210 49.24%
April 78,253 144,804 54.04%
May 99,861 173,403 57.59%
June 224,086 288,370 77.71%
July 248,412 308,815 80.44%
August 267,753 316,683 84.55%
Overall 970,663 1,338,285 72.53%
In Section 4.2, we discuss the block floor profit for builders. Like the two sides of the same coin, we now focus on rollups. Rollups are incapable of determining the exact number of type-3 transactions that builders will choose to include in a single block. Furthermore, rollups also do not know the pricing from other rollups. Consequently, we present the price for type-3 transactions in equilibrium.
Type-3 Transaction Price.  In equilibrium, for any type-3 transaction $b_{k}\in\mathbb{B}$, the expression for the price is
(3) $p_{k}^{\prime}=F_{tx}(b_{k}).$
For any given rollup, selecting a price below $F_{tx}(b_{k})$ will increase the likelihood that its type-3 transactions will not be chosen, thus diminishing the probability of verifying Data Availability. Conversely, selecting a price above $p_{k}^{\prime}$ will incur additional costs. For rollups, since the profit of type-2 transactions in the public mempool is not fixed, giving $p_{k}^{\prime}$ is sufficient to ensure that the builder will definitely choose $b_{k}$. The reason is that the profit of the builder after selecting this transaction will be greater than other remaining type-2 transactions under the same size.
Efficient Type-3 Transaction.  The efficient type-3 transactions here are that the specified profit of builders $p_{k}$ slightly exceeds the equilibrium price $p_{k}^{\prime}$. A threshold for this measurement has been established at an increase of 20% in $p_{k}^{\prime}$. Table 2 provides monthly efficient type-3 transactions. there are a total of 367,622 efficient transactions after EIP-4844, which is only 27% out of the whole. What’s more, from the delay perspective, efficient type-3 transactions with lower costs without significantly increasing delay. Based on our data, their average delay is only 0.78 seconds longer than inefficient type-3 transactions. Compared to Ethereum’s 12 seconds block interval, this difference is very small. We find that the proportion of efficient transactions is getting lower as time goes by, which means that rollups are becoming less sensible in their bids for type-3 transactions. We delve into the reasons for rollup’s reduced efficiency in Appendix F, which may be due to the fact that the volume of transactions in public mempool has been shrinking, with more than 10 million fewer transactions in August than in March since EIP-4844.
5.3. Flawed Strategies and Other Usage
[…]
Figure 7. The behaviors and amount of flawed strategies.
In addition to inefficient transactions, we also identify two existing flawed behaviors, Continuous Sending and Double Sending, and another usage, blobscription as shown in Figure 7. A Continuous Sending pattern is when the type-3 transaction sender sends more than one transaction in succession that are ultimately contained within the same block. These blobs can be put into a single type-3 transaction instead of multiple ones, which would require duplicate payments. Double Sending occurs when a rollup transmits the identical blob multiple times. A type-3 transaction that carries a second identical blob is invalid because the first successful transaction completes the validation of the blob data. These two strategies account for 7% and 10% of the total transactions, respectively. Additionally, it is important to mention that blobscription is observed in 106,051 transactions in the network, representing 8% of the total in general. This section evaluates the impact and economic loss to the network of the flawed strategies, showing the loss of 186.92 ETH and an average delay in network inclusion of 19 seconds.
5.3.1. Financial Loss.
[…]
Figure 8. The cumulative loss of flawed strategies.
We calculate the transaction fee (including base fee and priority fee) for duplicate transactions in Continuous Sending and Double Sending.  Figure 8 plots the cumulative losses from the Double Sending and Continuous Sending strategies for rollups after EIP-4844. The left y-axis shows the cumulative loss for these two strategies, and the right y-axis shows the total number of type-3 transactions. We use light blue bars to show the most rollups per month using these two strategies, and the yellow line represents cumulative losses. Since EIP-4844, rollups have lost 186.92 ETH. while Scroll sends the largest volume of transactions using the two inefficient strategies, it is worth noting that the number of these strategies has been decreasing since May, with less than 1 transaction in August. Another finding is that the use of the continuous sending strategy almost stopped in April, this is because Linea sent significantly fewer type-3 transactions in April.
5.3.2. Network Delay.
The presence of a large number of blobscription transactions violates the original intent of the EIP-4844 design and interferes with the normal rollup validation process. Since a block can only carry a maximum of 6 blobs, the massive number of blobscription transactions in the public mempool may cause a delay in the normal type-3 transactions. We further measure the impact of blobscription on network latency. We analyze 100,504 blobscription transaction during a period of burst of blobscription usage. During this period, there were 151,688 type-3 transaction sent from the rollups, and 21,282 of them were affected by blobscription. The results show that the average latency of the affected type-3 transaction was 43 seconds, while the other type-3 transactions only had to wait 24 seconds. It is clear that the presence of blobscription in the network introduces a delay of nearly 19 seconds to the rollups. Our finding exposes that if a large number of blobs that are not used for data validation are consistently present within the network, this causes a huge impact on the process of uploading type-3 transactions for rollups, which in turn affects the efficiency of the rollup application.
6. Related Work
Ethereum Upgrades.  Literatures have explored several upgrades related to Ethereum, including the appearance of ERC-20 tokens (Wu et al., 2023a; Zhao et al., 2021; Lin et al., 2024; Wu et al., 2023b; Kong et al., 2023), emergence of NFT (Huang et al., 2024; Ito et al., 2024; Xiao et al., 2024), the London hard fork (Liu et al., 2022; Leonardos et al., 2023), and the transition from PoW to PoS (Heimbach et al., 2023b; Silva et al., 2020; Chi et al., 2024; McLaughlin et al., 2023). After these updates, detection frameworks have been proposed to address the security issues posed by cryptocurrency tokens (Victor and Weintraud, 2021; Li et al., 2022). Hu et al. (Hu et al., 2023) developed a model to detect Ethereum fraud, while Zhou et al. (Zhou et al., 2024) examined the use of NFT for arbitrage and airdrop acquisition and revealed that it is driven by a small number of dominant players. Regarding the Ethereum merge event, Heimbach et al. (Heimbach et al., 2023b) quantified and analyzed hard-fork arbitrage in lending protocols. The tokens transferred by arbitrageurs to centralized exchanges were valued at more than \$13 million, and were actually withdrawn from the platform lenders during the Merge of Ethereum. This study complements the analysis of EIP-4844 and provides new observations on network efficiency in the interaction between builders and rollups.
Builder Behaviors.  Several previous works, such as (Öz et al., 2023; Ovezik et al., 2024; Gupta et al., 2023), investigated builder behaviors in PBS. The landscape exhibited that builders had a significant trend of centralization (Heimbach et al., 2023a). Likewise, Wahrstatter et al. (Wahrstätter et al., 2023) elucidated how vertical integration could potentially extract more value and further centralize block construction. Öz et al. (Öz et al., 2024) illustrated the relationship between the profit and market share, and how private order flow intensified the centralization of the builder market. In addition, Yang et al. (Yang et al., 2024) focused on competition and efficiency in MEV-Boost auctions of builders, and Heimbach et al. (Heimbach et al., 2024) quantified the non-atomic arbitrage utilized by builders to profit in the block-building process. Aside from this, blockchain censorship was defined to measure the censorship of builders in (Wahrstätter et al., 2024; Wang et al., 2023), demonstrating neutrality and security pitfalls. However, none explored builder strategies from the perspective of block sizes and building choices. To the best of our knowledge, we are the first to measure builders’ strategies towards distinct transactions and evaluate their efficiency after EIP-4844, offering crucial insights for builders to refine their block construction.
Rollup Analysis.  Extensive studies on the performance of the rollup ecosystem have revealed crucial insights into its present dynamics and potential future advances (Palakkal et al., 2024; Chaliasos et al., 2024). The process of MEV extraction within Layer-2 rollups has been analyzed by Torres et al. in (Torres et al., 2024), while Avil et al. (Khalil and Rahman, 2024) investigated the arbitrage of NFT that occurs in rollups. Gogol et al. (Gogol et al., 2024b) examine the effects of inscriptions on rollups. They find that the rise in meme transactions leads to reduced gas fees for rollups. Furthermore, Stefanos et al. (Chaliasos et al., 2024) provided insight into zero-knowledge rollups and developed a structured approach for their evaluation. The research presented in (Crapis et al., 2023) introduced an economic framework to examine EIP-4844 and its impact on rollups for reducing the cost of mainnet data. Our research provide the first rollup behavior evaluation and offers our insights on efficient strategies, which is useful for rollups to optimize their strategies in the post-EIP-4844 era.
7. Conclusion and Future Work
In this paper, we find that the relationship between builder and rollup is like the two sides of the same coin. In the one side, despite builders increase the overall block size, they still have more than 29% of blocks are inefficient. In the other side, a considerable portion of type-3 transactions from rollups paying unnecessary fees. Overall, our findings shed light on the challenges and opportunities that EIP-4844 has created for both builders and rollups. Our work paves the way for more efficient strategies that can mutually benefit the builder and rollup ecosystems, driving the evolution of Web3 infrastructure towards a more scalable and economically viable future.
We outline potential directions to optimize the block building process and Proto-Danksharding. After EIP-4844, the block building problem is essentially a multidimensional knapsack-constrained optimization problem (Mikerah, 2023), and how to build a more optimal block remains a future problem to be solved. Second, it is proved that a block carrying 6 blobs has no significant effect on propagation (Wahrstätter, 2024b, a; ethPandaOps, 2024). Based on the current transaction volume, the builder’s optimal block strategy is to put all blobs into one block. However, as the rollup ecosystem grows, the number of blobs that can be carried in a block will increase, and the potential missing rate will increase. Therefore, it is necessary to investigate the optimal size of blocks with more blobs.
Appendix A Blob Packing Strategies
[…]
Figure 9. The landscape of type-3 transactions after EIP-4844. The y-axis represents the daily usage and different colors indicates various rollups.
In Figure 9, we plot the blob number distribution after EIP-4844. Rollups that send one blob per type-3 transaction dominate. Most rollups use relatively fixed strategies. In addition to that, 2-4 blobs per transaction is a minimal amount.
Appendix B Delay of Type-3 Transactions
In this section, we focus on the two factors that influence the inclusion delay for type-3 transactions.
[…]
The delay of different quartiles type-3 transactions.
Figure 10. The delay of different quartiles type-3 transactions.
Non type-3 transactions with more priority fee would wait for shorter inclusion delay. However, since the size of type-3 transactions is significantly larger and thus may influence on builder’s efficiency. In Figure 10, we analyze the efficiency of builders in the public mempool spanning from EIP-4844 in March 2023 to the end of June. We rank the priority fees from lowest to highest and analyze the mean daily latency for type-3 transactions in both the highest and lowest quartiles. We select transaction delays within the range of 95% to mitigate the impact of extreme values. We also remove the result from March 27 to April 2 for the introduction of Blobscriptions (Blobscriptions, 2024) which spike the blob fee market.
The blue line denotes the delay for transactions in the 25% quartile (attributing lower priority fees), while the red line indicates the delay for transactions in the 75% quartile (attributing higher priority fees). It was observed that the lesser the priority fee paid, the longer the transaction delay. The pale blue bar shows the monthly average difference in delay between these two sets of transactions. The delay disparity between transactions with lower priority fees and those with higher priority fees has progressively grown. The difference has increased nearly 5-fold, from 1.54 s to 7.00 s. This finding shows that builders are becoming more and more efficient as the market evolves.
[…]
Inclusion delay and blob number.
Figure 11. Inclusion delay and blob number.
Figure 11 plots the relationship between the blob number, which is the transaction size, and the average latency of the transactions after EIP-4844 and fits the trend in the inclusion delay. The y-axis represents the delay time and the x-axis represents the number of blobs per transaction. We use bars with graduated colors from blue to green to represent the number of transactions. It is worth noting that the average latency for transactions with 3 blobs is a staggering 1200 seconds. Note that we separate data before and after June, when the builder market becoming more and more efficient. It may illustrate that all builders have transitioned from being previously chaotic to being ready. In the period after June, it is very clear that as the transaction size increases, the inclusion delay of the transaction increases correspondingly, a phenomenon that can be seen in the fitted folded line. This trend suggests that, as all block builders take block size into account, meaning that their algorithms are becoming efficient. This conclusion proves our result in Section 4.3.
Appendix C Greedy Algorithm
Input: The transactions remaining in the public mempool $\mathbb{T}_{r}$ , the number of blobs $\beta$;
1
2for $tx_{j}$ in $\mathbb{T}_{r}$ do
3       $\text{ratio}\leftarrow\text{$v_{j}$}/\text{$s_{j}$}$;
4
5 end for
[…]
7
[…]
11 for $tx_{j}$ in $\mathbb{T}_{r}$ do
12       if $\mathrm{total\_size+\text{$s_{j}$}}\leq\beta\times 128$ then
13             $\mathrm{total\_size}\leftarrow\mathrm{total\_size}+\text{$s_{j}$}$;
14             $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{j}$}$;
15             $\mathrm{selected\_list}.\text{append}(\text{$tx_{j}$})$;
16
17       end if
18
19 end for
[…]
22
Algorithm 1 Greedy Algorithm
We use a greedy algorithm as shown in Algorithm 1, considering the size and priority fee per transaction, calculating the priority fee ratio for each transaction and sorting by the ratio before processing.
Appendix D Invisible Profits for Builders
[…]
The profit of type-3 transactions size in the public mempool among builders.
Figure 12. The profit of type-3 transactions size in the public mempool among builders. Potential profits that can be extracted from the public mempool exceed those from type-3 transactions.
In Figure 12, we plot the results for six months in different colors among the main builders. The right y-axis of Figure 12 shows the profits that builders can extract from the public mempool, corresponding to the blue bars, while the left y-axis represents the transactions that are actually extracted from type-3 transactions, corresponding to the green bars. It reveals that builder can extract more profit in public mempool, but instead they chose to include the much cheaper type-3 transaction into block. The cumulative profits of other transactions for each builder are overwhelming compared to type-3 transactions, which explains that type-3 transactions give significantly lower priority fees than the other transactions, and do not compensate at all for the impact of their size. In addition, we observe that rsync builder may choose fewer type-3 transactions in their blocks because Vanilla builders have a lower market share but are more profitable from both type-3 transactions and other transactions from public mempool.
[…]
Figure 13. The relationship between delays and maxPriorityFeePerGas among rollups. Similar price from rollups have uneven inclusion latency.
Figure 13 illustrates the average delay and average priority fee of different rollups in August. The y- and x-axes indicate the average latency and priority fee, respectively, while each point denotes a particular rollup. In the context of the builder sorting algorithm being absolutely efficient, the inclusion delay of type-3 transactions with the same priority fee should be close. However, in Figure 13, rollups that offer close to average priority fees have completely different inclusion delays. For example, Optimism has a higher priority fee than Zircuit, Kroma, and Arbitrum, but the latter still has a shorter waiting time for type-3 transactions than Optimism. On top of that, Kroma and Arbitrum, etc. have transaction latencies that are very comparable to Rari, but their transaction priority fees are just half that of Rari.
Appendix E Private Channels between Builder and Rollups.
We compare on-chain data with data recorded by public mempool and label data that bypassed public mempool as private transactions from EIP-4844 to August. Our final private type-3 transaction dataset consists of 3,153 private transactions, of which 826 transactions come from a rollup called Lisk, which accounts for 98% of the type-3 transactions issued by this roll-up. Interestingly, 1,405 of these 1,484 private transactions from Taiko were built into the block by the Titan builder to be on the main net, and 795 transactions from Lisk also built into Titan’s blocks.
Table 3. The Top 5 private type-3 transaction senders
Address Rollup Counts
0x000000……663bcbe0 Taiko 1484
0xa6Ea2f……9cd6fe24 Lisk 826
0x2C169D……fe260ca7 StarkNet 179
0x0D3250……f790ec99 zkSync 128
0xcF2898……0b4cfc8f Scroll 147
Appendix F Reduction in Public Mempool
[…]
Figure 14. The inclusion delay for type-3 and non type-3 transactions over time.
A smaller public transaction block size may lead to an increased transaction inclusion delay. To evaluate it, we display the overall transaction latency in Figure 14. Nevertheless, transactions in the public mempool did not experience extended inclusion delays, and the overall trend consistently decreased. The decrease in public mempool transactions explains why there hasn’t been an increase in latency. In fact, there were 28,012,234 transactions in the public mempool in August, while in March, there were 41,500,537 transactions, highlighting a decrease of 13,488,303 transactions.
Table 4. Granger causality test of block size and mempool size
$lag_{x}=lag_{y}$ F-value p-value
1 35011.6924 0.00
2 26616.7382 0.00
3 20894.1280 0.00
4 17072.5844 0.00
5 14266.8419 0.00
We use Pearson’s correlation coefficient and the Granger causality test to further evidence the correlation between the size builder extracts from the public mempool and the quantity of the public mempool. The Pearson correlation coefficient is 0.4335 with a p-value of 0.00. The p-value indicates the relationship between the overall size of transactions extracted by the builder from the mempool and the decreasing quantity of mempool. We also perform a Granger causality test in Table 4, indicating that a decrease in size correlates with a reduction in the number of transactions in the public mempool. The reduction in the public mempool could be associated with EIP-4844. This is because more transaction senders are choosing to use the lower-priced Layer-2 after EIP-4844 instead of the Ethereum mainnet (Pokorny, 2024).