### v3

# The Price of Cheaper Data: Measuring the Strategic Inefficiencies in the Post-EIP-4844 Ethereum Market

Abstract.
High transaction fees on Ethereum have long hindered its scalability and user adoption. The recent Ethereum EIP-4844 upgrade aims to alleviate the scalability issue by introducing the “blob”, a new data structure for Layer-2 rollups. Instead of using expensive blockchain storage, blobs provide a cheaper, separate data layer with its own fee market, which drastically lowers data availability costs. This change, while lowering transaction costs, has created a new high-stakes economic game for block builders and rollups. However, the dynamics of this game remain poorly understood. In this paper, we conduct the first large-scale empirical analysis of the post-EIP-4844 ecosystem, leveraging a dataset of $319.5$ million transactions, including $1.3$ million blob-carrying type-3 transactions. Our analysis demonstrates clear evidence of a structural shift towards the utilization of cheap blobs over expensive transactions for Layer-2 data posting: while average block size grew $2.5\times$, the space consumed by expensive transactions in the public mempool shrank from over 150 KB to just 30 KB. Yet, this scalability success masks widespread economic inefficiencies on both sides of the market. On the builder side, $29.48\%$ of blocks containing blobs are constructed sub-optimally, earning less revenue than that could be achieved otherwise. On the rollup side, we identify that flawed submission strategies have led to $186.92$ ETH in direct losses and average inclusion delays of over $19$ seconds. Moving beyond characterization, our work offers actionable solutions. For builders, we develop an optimal pricing model derived from a formal first-price auction framework, allowing builders to make provably profitable inclusion decisions. For rollups, we prove that batching multiple blobs into a single transaction is a dominant, utility-maximizing strategy. Our work presents the first data-driven characterization of the post-EIP4844 market, revealing critical insights for the future development of Ethereum’s scalability.
Ethereum, Layer-2, Builder, Rollup, EIP-4844, Blob
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: General and reference Empirical studiesccs: General and reference Measurementccs: General and reference Evaluation
1. Introduction
1.1. Background
Ethereum is a decentralized, blockchain-based smart contract platform. The ability to execute code trustlessly allows Ethereum to serve as a foundational layer for a wide array of decentralized applications, including decentralized Finance (DeFi), decentralized voting systems, and stablecoins. Executing transactions and smart contracts on the network consumes computational resources, which requires a fee to manage demand and incentivize network participants. This fee is calculated using“Gas”—a unit that measures the complexity of an operation—and is paid in the platform’s native cryptocurrency, Ether (ETH). This economic model has proven highly effective, fueling the platform’s widespread adoption. Data from Glassnode indicates that Ethereum has become one of the world’s leading blockchains, with its total address count reaching $250$ million and surpassing Bitcoin in several key on-chain activity metrics (Glassnode, 2025).
As the Ethereum user base has grown, the mainnet has faced significant network congestion, leading to limited throughput, extended transaction confirmation times, and prohibitively high fees. To address these critical scalability limitations, the community has championed external solutions known as Layer-2 protocols. Among the Layer-2 protocols, Rollups have emerged as a prominent approach. Rollup solutions enhance efficiency by bundling hundreds of transactions into a single batch on a separate chain and then submitting a compressed summary of this batch to the Ethereum mainnet for final settlement. By offloading the bulk of computation and execution in this manner, Rollups drastically reduce congestion and lower costs for users, all while inheriting the robust security guarantees of the mainnet.
The growing adoption of Layer-2 solutions, particularly Rollups, placed significant economic pressure on their operators. To ensure data availability and security, these protocols were required to post compressed transaction data to the expensive “Calldata” space on the Ethereum mainnet—used for verifying data in Layer-2 transactions As user activity scaled, the frequency and cost of these “Calldata” submissions increased commensurately, creating a major operational burden. At its peak, this “data tax” on Calldata often cost major Rollups a collective sum exceeding $10,000$ USD per day, according to analyses on Dune Analytics (0xRob, 2025). To alleviate this, the Ethereum Dencun upgrade introduced EIP-4844, which created a dedicated and more cost-effective data space for Rollups known as “blobs”. Accessed via a new Type-3 transaction, blobs allow Rollups to offload data from Calldata, thereby significantly reducing their operational costs (Ethereum.org, 2024a; Buterin et al., 2023). This new dynamic structurally alters the decision-making for both block builders and rollup operators. Block builders, whose primary task is to construct the most profitable block by sorting transactions (Daian et al., 2020; Heimbach and Wattenhofer, 2022), now face a critical dilemma. Including large, blob-carrying transactions may increase fee revenue, but it also increases the block’s propagation time, heightening the risk of it being orphaned and losing all profit. Concurrently, rollup operators must compete to have their Type-3 transactions included by these discerning builders. This has led to multifaceted strategies (Crapis et al., 2024), where Rollups must decide on optimal blob bundling and offer compelling priority fees to incentivize builders to accept the added risk. The result is a new, complex marketplace for blob inclusion with diverse bidding behaviors (Cui, 2024).
Failing to properly understand this market can have adverse consequences. For builders, a persistent reliance on suboptimal block construction strategies may prevent them from capturing the revenue opportunities provided by type-3 transactions. Over time, this revenue gap erodes their competitiveness relative to faster-adapting peers, ultimately forcing them out of the builder market (Wang et al., 2025). For rollups, inefficient blob-carrying strategies result in consistently higher inclusion costs and longer delays, a burden that is inevitably passed on to the users of their Layer-2 networks.
1.2. Contribution
Our research is driven by the following three research questions, where efficiency is understood in terms of profitability.
  • RQ1: What are current builder strategies for block construction, and are they efficient?
  • RQ2: What are current rollup strategies for submitting blobs, and are they efficient?
  • RQ3: What is the profit-optimal strategy for builders and rollups?
To answer the questions and illuminate the unexplored dynamics of the post-EIP-4844 market, we begin by constructing the first large-scale dataset dedicated to analyzing the strategic interactions between builders and rollups in this new era. This comprehensive dataset, encompassing over $319$ million transactions—including $1.3$ million Type-3 transactions, serves as the bedrock for our entire investigation. We have made the dataset publicly available to benefit the broader research community. [^1] We would make our dataset public in the published version
Our investigation first provides clear evidence of the profound structural shift EIP-4844 induced in the market. We observed that as Layer-2s moved their data posting from expensive calldata to cheap blobs, the average block size grew $2.5\times$ to nearly $400$ KB. Simultaneously, the space consumed by traditional transactions from the public mempool shrank dramatically, from over $150$ KB to just $30$ KB. This confirms that builders are actively reallocating block space to accommodate blobs. The fundamental change in builder strategy confirmed their adaptation to the new market. This, in turn, motivated our investigation into its economic efficiency.
However, our investigation reveals that this strategic shift led to widespread economic inefficiency. We identify a significant “crowding out” effect, where builders frequently accepted blob transactions even when their revenue failed to cover the opportunity cost of excluding more profitable transactions readily available in the mempool. This inefficiency was particularly stark in March 2025 (the month following the upgrade), where potential profits from the mempool were over five times greater than the revenue builders actually gained from blobs. Across our entire observation period, we quantify this issue, finding that a striking $29.48\%$ of all blocks containing blobs were unprofitable for this very reason.
The situation for rollups proved to be equally concerning. Our analysis uncovers that their post-EIP-4844 strategies are far from optimal, leading to both direct economic losses from flawed strategies and significant network externalities from blob misuse. Specifically, we identify two prevalent inefficient submission practices. The first, Continuous Sending (accounting for $7\%$ of transactions), occurs when rollups pay redundant fees by sending multiple transactions that could have been bundled into one. The second, Double Sending ($10\%$ of transactions), involves wasting fees on submitting identical, useless blobs after the first one has already been validated. Combined, these self-inflicted inefficiencies have cost rollups a quantifiable loss of at least 186.92 ETH. Furthermore, rollups suffer from the misuse of blob space by external actors. We identify the rise of “Blobscriptions”—a practice of using blobs for on-chain inscriptions rather than their intended purpose of rollup data validation. This creates a negative externality by consuming scarce resources (a maximum of six blobs per block) and congesting the network for legitimate rollups. The impact is significant: during periods of peak activity, this congestion imposed an additional inclusion delay of nearly 19 seconds on legitimate rollup transactions—increasing their average wait time from 24 to 43 seconds—and severely degrading overall service performance.
To address these widespread inefficiencies, we then propose formal solutions. We construct a game-theoretic model to capture the complex interplay between builders and rollups, and from this model, we derive actionable, optimal strategies to correct the identified market failures. For builders, we propose a new pricing framework to ensure provably profitable block construction. Our results also show that, as the number of type-3 transactions increases, the price obtained by builders rises and gradually converges to the revenue that rollups derive from these transactions. For rollups, we find that blockspace constraints do not significantly affect the selection probability of large type-3 transactions. Under these conditions, we then formally prove that bundling multiple blobs into a single transaction is a dominant, cost-saving strategy that maximizes their utility. In essence, our work not only diagnoses the nascent problems of the EIP-4844 market but also provides a clear, evidence-based roadmap toward a more efficient ecosystem.
To summarize, this paper presents the first systematic analysis of the economic dynamics and market strategies on Ethereum following the EIP-4844 upgrade. Our main contributions are:
  1. (1)
    We construct a large-scale dataset to analyze the builder and rollup market in the post-EIP-4844 era. Based on that, our analysis reveals that EIP-4844 successfully altered Ethereum’s transaction structure and significantly reduced data fees.
  2. (2)
    We conduct a detailed empirical study of builder and rollup strategies, revealing that both parties operate sub-optimally. Our findings show that builders use inefficient pricing, while rollups exhibit flawed behaviors like redundant blob submissions, resulting in missed revenue and unnecessary costs.
  3. (3)
    We develop an optimal pricing model to enable builders to make provably profitable decisions. For rollups, we formally prove that bundling multiple blobs into a single transaction is a dominant, cost-saving strategy.
1.3. Organization
This paper is structured as follows. Section 2 introduces the preliminaries of Layer-2 Rollups, Ethereum Builders, and the details of the EIP-4844 upgrade. Section 3 describes our dataset creation process. Section 4 examines builder strategies in the post-EIP-4844 era by analyzing block composition, revenue, and efficiency, identifying prevalent inefficiencies in their current behavior. Section 5 then evaluates rollup behavior regarding blob and type-3 transaction usage, highlighting their inefficient pricing strategies and the resulting implications. Section 6 develops a game-theoretic model between builders and rollups to derive optimal pricing strategies for builders and utility-maximizing strategies for rollups. Finally, Section 7 reviews related work, and Section 8 concludes the paper.
2. Preliminaries
[…]
Illustration of the block building process in Ethereum.
Figure 1. Illustration of the block building process in Ethereum. Builder is divided into MEV-Boost builders and Vanilla builders according to their transaction source. Their block would contain type-3 transactions introduced by EIP-4844.
In this section, we establish the foundational knowledge required to analyze the post-EIP-4844 market. We begin by defining the operational mechanics of key actors, namely builders and rollups. We then delve into the specifics of the EIP-4844 proposal, explaining how data blobs and Type-3 transactions function. A clear grasp of these components is crucial for understanding the subsequent analysis presented in this paper.
2.1. Layer-2 Rollup
Rollups are Layer-2 scaling solutions designed to improve the efficiency and throughput of the Ethereum network by offloading transaction processing from the Ethereum mainnet (Ethereum.org, 2024b). To protect the data availability of Ethereum (Pomerantz, 2024), rollups work by bundling multiple transactions into a single batch. After bundling, the rollup smart contract on Ethereum updates the state with a single, compressed transaction that includes the aggregated data of all individual transactions. This compressed transaction will be submitted to the mainnet, thereby reducing the computational load and gas fee for each individual transaction. Rollups maintain the security and decentralization of Ethereum while providing faster and cheaper transactions, making them a crucial component of the Ethereum scalability roadmap (Smith, 2024).
Rollups come in two main types: optimistic rollups (Guimarães, 2024) and zero-knowledge rollups (Wackerow, 2024b). Optimistic rollups assume transactions are valid and only run fraud proofs when discrepancies are detected, whereas zero-knowledge rollups use cryptographic proofs to validate the correctness of the bundled transactions. Both solutions have unique trade-offs and use cases, but share the common goal of making Ethereum more scalable without sacrificing security.
2.2. Ethereum Builder
With an upgrade known as the Merge (Foundation., 2023) in 2022, Ethereum has transitioned from a PoW consensus, which is energy intensive, to a PoS mechanism. The Proposer Builder Separation (PBS) mechanism has been proposed as an important development after the Merge (Ethereum.org, 2023). The traditional role of validators, formerly known as miners, is divided into separate entities: MEV-Boost builders and proposers (Buterin, 2021a, b). MEV-Boost builders are responsible for block construction. They select transactions from public mempool (Wackerow, 2024a) or private channels (Wang et al., 2025). The constructed blocks are then forwarded to proposers, accompanied by fee commitments offered by the builders. Proposers endorse the block with the highest fee for the finalization of the blockchain.
Currently, it is optional for Ethereum validators to participate in PBS (Wahrstätter, 2024). Validators who opt not to use PBS and instead build blocks independently are known as Vanilla builders. Vanilla builders receive no private transactions because they are not specialized in block building tasks. They complete both block building and validation tasks simultaneously and currently account for 10% of all builders in the network (Wahrstätter, 2023).
Figure 1 illustrates the current architecture of Ethereum. MEV-boost builders receive transactions from both the public mempool and private channels, whereas Vanilla builders rely solely on the public mempool. Despite differing transaction intake mechanisms, both types of builders contribute to the same outcome — a competitive blockspace market in which users bid for faster inclusion by offering higher priority fees (Liu et al., 2022). Users seeking expedited confirmation increase their priority fees, while those with lower urgency accept longer delays by paying only the base fee.
2.3. EIP-4844 & Type-3 Transaction
Rollups use calldata to store state roots and proofs (Pomerantz, 2024), ensuring that the rollup state could be re-established even if a transaction was challenged. Calldata is the payload contained in a transaction that can be used to store arbitrary information on the chain. While calldata is cost-effective for smaller amounts of data, it is permanently stored in Ethereum state and is therefore relatively expensive for larger data requirements. The high cost of calldata storage becomes a bottleneck for scaling.
In order to improve the scalability and data availability of the network, Ethereum proposed a new sharding design, Danksharding (Ethereum.org, 2024b). Danksharding is a unified sharding design that differs from traditional sharding by managing data storage and availability in a more streamlined way. Instead of increasing the space for transactions, Ethereum sharding provides more space for blobs of data, which the Ethereum protocol itself does not interpret.
EIP-4844, also known as Proto-Danksharding, implements the same transaction formats and validation rules. EIP-4844 was launched in March as a core part of the Dencun upgrade in 2024 (Ethereum.org, 2024a), thereby facilitating the network’s preparation for full Danksharding by allowing the inclusion of large data in transactions (Pomerantz, 2024). The EIP-4844 represents a significant milestone in Ethereum’s roadmap (Ethereum.org, 2024c), aiming to improve the scalability, security, and usability of the network.
There were three types of transaction before EIP-4844: type-0, type-1 and type-2 transaction. Note, type-2 transactions always bring a priority fee for inclusion by builders quickly. EIP-4844 implements on a new type: type-3 transactions (Park et al., 2024). These are similar to regular type-2 transactions, but carry an additional data component called a ‘blob’. A blob is a large, temporary data payload attached to a type-3 transaction. Unlike calldata, blob data is not stored permanently on-chain. Instead, only a commitment to the blob is included on-chain, significantly reducing storage costs (Pomerantz, 2024). These oversized blobs (128 KB) are much cheaper than the equivalent amount of calldata because they are not permanently stored on Ethereum (Pomerantz, 2024), and rollups can also send their type-3 transactions from public mempool or private channels. This new type of transaction meets the growing need for efficient data processing. The blob data remains inaccessible during the execution of the Ethereum virtual machine, and the validator can only see the commitment to the blob, but these blobs can be used for Layer-2 scaling solutions and other off-chain protocols that require large amounts of data throughput (Gogol et al., 2024a).
3. Data Collection Methodology
We generate our data in two stages to study the dynamic builder and rollup strategies. First, to support the analysis, we gather public on-chain data, including block, transaction, and receipt data. To collect transactions in the public mempool, we use archived data from Flashbots (Flashbots, 2023). We decode the data that is used to identify block builders to label both MEV-Boost builders and Vanilla builders, and collect rollup address to identify rollup applications and others. As far as we are aware, this is the first time such an extensive dataset with labels has been constructed. To facilitate further research, our labeling dataset would be released in the published version.
3.1. Data Collection
On-chain Data.  To build the on-chain dataset, we deploy the Ethereum execution client Erigon (Ledgerwatch, 2024) and the consensus client Lighthouse (Prime, 2024). We collect on-chain data covering the period from December 1, 2023 to August 31, 2024. The on-chain dataset contains 1,967,415 blocks, 319,529,950 transactions and receipts. The dataset also includes the hash address of each blob in 1,336,822 type-3 transactions.
Public Mempool Data.  We next acquire the public mempool archive data from Flashbots Mempool Dumpster (Flashbots, 2023), which is updated every day. The mempool data provides the transaction inclusion delay and the information such as the type and size of a transaction (Hager, 2022). Note that the transaction inclusion delay is calculated from the time interval between the first time the transaction is detected and the time when it is included in a block. The public mempool data range from the same period of on-chain dataset and have 318,175,240 public transactions.
3.2. Labeling
Builder Labeling.  To identify the builder of a particular block, the usual approach is to first check if it is a vanilla block and then query the endpoints from the relay. However, this is complex and time consuming. To better match the builder and its blocks, we decode the extra_data from the hexadecimal string to the origin value. To measure our coverage, we compare our builder labeling dataset with existing public dataset (Network, 2023) created using builders’ public keys. This confirms that our dataset covers over 300 vanilla builders which were ignored in other datasets. Our labels contain 56 MEV-Boost builders and 318 Vanilla builders.
Rollup Labeling.  To label the observed rollups in Layer-2, we collect the official on-chain address disclosed by (hildobby, 2024; Tran, 2024) to match the corresponding rollups. However, we find that not all rollups are correctly identified. To remedy this, we note that some rollups add their chain ID to the “to” address in their transactions. Thus, to label more rollups, we extract the chain ID of these rollups and request the chain ID using the Chainlist API [^2] https://chainlist.org/). The labels we gather cover 35 main rollups.
To ensure that our dataset is accurate, we compare each address with their official documents (Arbitrum, 2025; Taiko, 2025; Scroll, 2025). To our knowledge, our rollup labels offer by far the most coverage. Additionally, we also find that a large number of type-3 transactions are used for blobscriptions (Blobscriptions, 2024), to filter these transactions, we use the Ethscription API to collect these blobscriptions transactions. [^3] https://api.ethscriptions.com/v2/ethscriptions/ Utilizing this API, it is possible to retrieve a blobscription via its transaction hash. We identify and label all instances within type-3 transactions belonging to blobscriptions. Consequently, our labeling process encompasses 106,051 type-3 transactions used for blobscriptions. We identify all unlabeled addresses in our type-3 transaction dataset and annotate 5,655 Blobscription addresses following EIP-4844.
4. Characterizing Builder Behaviors
To analyze how builders select transactions and whether they are efficient after EIP-4844, as stated in RQ1 in Section 1, we first quantify block-building strategies following the introduction of large-volume type-3 transactions. We then propose a model to evaluate builders’ efficiency—interpreted throughout this work in terms of profitability—and analyze whether current market behavior aligns with this notion.
4.1. Block Building Strategies
(a) Block size.
[…]
Block size.
(b) Size excluding type-3 tx.
[…]
Type-3 excluded block size.
(c) Non type-3 size from mempool.
[…]
Block size from public mempool.
Figure 2. The average size in block among builders.
[…]
Average block size from builders.
To demonstrate the difference of builder’s strategies towards block construction after EIP-4844, we analyze the block size alongside changes in specific components. In 2(c), we present the average block size, the size of non type-3 transactions within the block, and the size of non type-3 transactions from the public mempool within the block, respectively. These cover the period from December 2023 to the end of August 2024. First, in 2(a), the black line represents the overall average block size, which is generally on an upward trend. EIP-4844 leads builder to switch their strategy, and the total size of the block increases significantly, from less than 150 KB to near 400 KB (roughly a 2.5-fold increase). Second, the blue and green lines represent the Flashbots builder and the rsync builder, respectively, where Flashbots changed its strategy in May, choosing to exclude type-3 transactions from the block, and then continuing to include them in July.
To measure the builder’s varying behavior towards other non type-3 transactions after the upgrade, in 2(b) we observe the changes in the size of the non type-3 transaction portion of the block (this is the space occupied by all non type-3 transactions within a block). Clearly, after exceeding 300 KB in January, the average block size without type-3 transactions began to decline to 80 KB by June, reducing by almost four times. This indicates the efficacy of EIP-4844. The yellow and orange lines represent the strategy of the Beaverbuild builder and Vanilla builders, who were the first to implement the strategy of decreasing the total size of other transactions within a block following EIP-4844. Nearly 60% of the market share is held by these builders, and the majority of the blocks they constructed are below the average size prior to May. By March, Vanilla builders transitioned to creating smaller blocks of other transactions, possibly due to not receiving a higher volume of orders and lacking the professional expertise to build blocks. Notably, starting from June, the top five market share builders, including Flashbots builder and Titan builder who previously built the largest blocks, have begun creating smaller sized blocks.
In conjunction with 2(a), this finding suggests that the builders are reducing the size of the other parts in order to include the type-3 transaction in the block. It is worth noting that the volume size reduces to below 80 KB after June, and remains stable at around 80 KB thereafter. That is, the increase in type-3 transactions has squeezed space for type-2 transactions.
2(b) depicts the non type-3 transaction sizes within a block. These can be categorized into two channels: public mempool and private channels. Due to the exclusive access of the private channel to specific builders, in 2(c), we illustrate the size of other transactions from the public mempool within a block to further analyze the strategy behind the smaller size. The general trend observed in 2(c) is largely similar to that of 2(b). This suggests that the reduction in block size is due to builders minimizing the size of transactions taken from the public mempool. It is concerning to recognize that all builders employ a detrimental tactic to select small-volume transactions from the public mempool, potentially causing extended delays for numerous transactions that should have been included in the block.
4.2. Efficient Strategy for Builders
(a) The profit of type-3 transactions and public mempool over time. The left y-axis shows the potential profit and right y-axis represents the profit difference.
[…]
(b) The proportion of inefficient blocks per builder. The number above the bar is the total blocks specific builder built.
[…]
Figure 3. The profit distribution and efficient block. On left is the different profit distribution from type-3 transactions and public mempool. On right is the efficient proportion of different builders.
To answer the RQ1, we derive the efficient strategy for builders in this section. For builders, the decision between type-2 and type-3 transactions constitutes a significant consideration. Both transaction types provide builders with profits via the ‘maxPriorityFeePerGas’ field. For brevity, in this paper, we substitute the term ‘transaction price’ for it. The previous section has shown that builders reduce the volume of type-2 transactions and choose to select more type-3 transactions. To identify the most efficient strategy for deciding this ratio, we derive the model of the floor price of block space by using type-2 transactions in the public mempool.
Let $\mathbb{M}=\{m_{1},m_{2},\ldots,m_{u}\}$ denote the set of type-2 transactions remained in the public mempool and $\mathbb{N}=\{n_{1},n_{2},\ldots,n_{k}\}$ denote the set of type-3 transactions within the time window of the block construction. For builders, the research shows that the factors they are concerned about are the price and the block size (Wu et al., 2024; Wahrstätter et al., 2023). Therefore, for each type-2 transaction $m_{i}$ in $\mathbb{M}$, we denote its size as $s_{i}$ and its price as $p_{i}$. Then the transaction $m_{i}$ can be further denoted as $m_{i}(s_{i},p_{i})$. For each type-3 transaction $n_{j}$ in $\mathbb{N}$, we denote the number of blobs included in it as $q_{j}$ and the price as $p_{j}$. As specified by EIP-4844, each blob has a size of 128 KB. Therefore, the type-3 transaction can be further represented as $n_{j}(128q_{j},p_{j})$.
Floor Price. The floor price for block space usage by type-3 transactions represents the maximum profit that can be achieved by selecting from $\mathbb{M}$ with the same size. In the rest of the paper, we refer to this as the “floor price”. Formally, the floor price $P(n_{j})$ can be defined as
(1) $P(n_{j})=\max\sum_{m_{i}\in\mathbb{M}}v_{i},$
[…]
Intuitively, it represents the maximum profit that builders can obtain from the remaining type-2 transactions in the public mempool to compensate for the consumption of block size by type-3 transactions.
Efficient Strategy for Builders.  The floor price represents the highest profit that builders could generate from type-2 transactions remaining in the public mempool, subject to the block space usage of type-3 transactions. So for builders, the efficient strategy is to choose type-3 transactions if the profit satisfy
(2) $p_{j}\geq P(n_{j}).$
And if Equation 2 does not hold, then the efficient strategy is to directly select type-2 transactions from the public mempool. Blocks wherein all type-3 transactions adhere to the efficient strategy are categorized as efficient blocks. We find that there are still many inefficient blocks. Specific data and analysis will be provided in the next section.
4.3. Efficiency of Builders
Input: The transactions remaining in the public mempool $\mathbb{T}_{r}$ , the number of blobs $q$;
1
2for $m_{i}$ in $\mathbb{T}_{r}$ do
3  $\text{ratio}\leftarrow\text{$v_{i}$}/\text{$s_{i}$}$;
4
5 end for
[…]
7
[…]
9 $\text{total\_tips}\leftarrow 0$;
10 $\text{selected\_list}\leftarrow[]$;
11 for $m_{i}$ in $\mathbb{T}_{r}$ do
12  if $\mathrm{total\_size+\text{$s_{i}$}}\leq q\times 128$ then
13     $\mathrm{total\_size}\leftarrow\mathrm{total\_size}+\text{$s_{i}$}$;
14     $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{i}$}$;
15     $\mathrm{selected\_list}.\text{append}(\text{$m_{i}$})$;
16
17  end if
18
19 end for
[…]
21 return $\text{total\_tips},\text{selected\_list}$;
22
Algorithm 1 Greedy Algorithm
Table 1. The overall inefficient block distribution over time.
March April May June July August Overall
Inefficient Block 26,450 41,779 47,351 31,540 27,965 23,139 198,224
Total Block 47,604 84,631 105,382 139,529 145,726 149,515 672,388
Inefficient Proportion 55.56% 49.37% 44.93% 22.60% 19.19% 15.48% 29.48%
To measure whether builders’ new strategies are efficient or not, we employ a greedy algorithm considering size and profit to measure whether the priority fee given to each blob can match the priority fee that can be extracted in the public mempool during the concurrent period. Note that before being executed, only the maximum priority fee, i.e. maxPriorityFeePerGas field, can be seen. This represents the profit for builders in one transaction. We use this value to calculate the potential profit for builders. To filter out transactions that builders have simulated as likely to fail, we filter out transactions in the public mempool with a delay of more than 100 seconds to make the calculation more accurate. We also provide the pseudocode of the greedy algorithm. The algorithm considers the size and priority fee per transaction, calculating the priority fee ratio for each transaction and sorting by the ratio before processing.
3(a) plots the profit that builders actually extract from type-3 transactions versus the profit that can be extracted from public mempool in the same period from March to August. We compute the profit difference between the profit of type-3 transactions and the profit that can be extracted from the public mempool, and indicate this value with red and yellow bars. The red bar indicates that this difference is negative, which suggests that the builders could have made more profit from the public mempool, but they chose to include the less profitable type-3 transactions in the block. Conversely, if the color is yellow, it indicates that the value is positive and the builder made more profit from type-3 transactions. It is clear that after EIP-4844, the builder misses significant value of the public mempool. As 3(a) shows, in March, builders can potentially make more than five times more profit from the public mempool as they can from type-3 transactions. However, this loss diminishes over time and reverses in July, when the builder makes more profit from type-3 transactions than it is able to make from the public mempool.
Table 1 illustrates the details from March to August, including the number of efficient and inefficient blocks and their proportion out of the 636,981 blocks containing type-3 transactions. We find that 196,090 blocks are inefficient, which is approximately 29.48%. Encouragingly, the percentage of inefficient blocks dropped from 55.56% immediately after EIP-4844 to 15.48% by August, indicating a trend toward a more efficient block building market. However, 15.48% remains a non-negligible share, underscoring that there is still considerable room for improvement.  3(b) shows the ratio of efficient blocks among major builders during the same period. The y-axis shows the proportion of efficient and inefficient blocks. Beaverbuild has a market share of almost 50%, and is the most efficient compared to other builders. In contrast, non-specialist builders, vanilla builders who do not specialize in block construction, have nearly 30% more inefficient blocks than beaverbuild, and more than half of the blocks they construct are inefficient.
In 4(a), we plot the results for six months in different colors among the main builders. The right y-axis of 4(a) shows the profits that builders can extract from the public mempool, corresponding to the blue bars, while the left y-axis represents the transactions that are actually extracted from type-3 transactions, corresponding to the green bars. It reveals that the builder can extract more profit in public mempool, but instead they chose to include the much cheaper type-3 transaction into the block. The cumulative profits of other transactions for each builder are overwhelming compared to type-3 transactions. This is because type-3 transactions give significantly lower priority fees than the other transactions, and do not compensate for the impact of their size. In addition, we observe that rsync builder may choose fewer type-3 transactions in their blocks and Vanilla builders have a lower market share but are more profitable from both type-3 transactions and other transactions from public mempool.
Answer to RQ1 EIP-4844 significantly altered block construction strategies: the overall block size expanded dramatically by approximately 2.5 times, growing from less than 150 KB to close to 400 KB. Concurrently, the available capacity for non type-3 transactions was reduced roughly fourfold from 300 KB to 80 KB, indicating that type-3 transactions crowded out other transactions. From March to August, 29.48% of the 636,981 blocks containing type-3 transactions were inefficient, with builders missing potential profits from the public mempool. In March alone, builders could have earned over 5× more from the mempool than from type-3 transactions. However, efficiency improved over time, with the share of inefficient blocks dropping from 55.56% post-EIP-4844 to 15.48% by August, driven largely by specialist builders like Beaverbuild, whose inefficiency rate is over 30% lower than that of non-specialist builders.
5. Understanding Rollup Strategies
After the overview in builder market, as outlined in RQ2 in Section 1, we now present a measurement of rollups to obtain a comprehensive understanding of their current strategies toward blob and type-3 transaction usage, as well as their efficiency. In this section, we examine the opposite side of the market, rollup’s strategy. We start by analyzing the different rollup behaviors after EIP-4844. We conduct an analysis about the efficiency of rollups and identify two inefficient practices and measure their impact.
5.1. Type-3 Transactions Usage
After the upgrade, in total 1,336,822 type-3 transactions have been sent to the Ethereum mainnet. 4(b) visualizes the daily share of type-3 transactions by the dominant rollups and a specific role after EIP-4844, in terms of the total number of blob-carrying transactions. We notice that the number of transactions for each rollup varies over time, with some rollups showing significant spikes or drops. One of the most noticeable spikes is in Blobscriptions, which began hitting the market at the end of March, reached its peak, and then began to decrease dramatically. The purpose of Blobscriptions is not to validate rollup data, but rather to publish it. Other rollups show varying increases and decreases in usage, which is related to the choices made by each rollup application. Obviously, the use of type-3 transactions has been increasing after EIP-4844, with more rollup applications using blobs to validate data. Instead of immediately using a blob to transfer data, some rollups (e.g., Taiko) only started adopting type-3 transactions in May of this year. However, there are still many rollup applications (such as Arbitrum and Base) that have been using type-3 transactions since EIP-4844.
(a) The profit of type-3 transactions size in the public mempool among builders.
[…]
(b) The landscape of type-3 transactions.
[…]
Figure 4. The landscape of type-3 transactions and blob usage.
In  5(a), we plot the blob packing strategy for each mainstream rollup from 13 March to 31 August. The different colors represent the number of blobs carried in type-3 transactions and the Y-axis represents the total number of transactions. It is clear that rollup makes different choices in selecting the number of blobs carried in a type-3 transaction. Of the 343,468 transactions issued by Taiko, it puts only one blob into the transaction, and the same strategy is used by Scroll, Starknet, and Zircuit. We further investigate the reasons for this in Section 6.2, and find that the latency of the transactions that carry only one blob is the lowest. Therefore, it is likely that this strategy is optimizing the latency of the rollup application so that the user’s transactions can be verified the fastest. In contrast, Base and Op Mainnet, which have been using blobs to send data, put as many blobs as possible into a transaction, and most of their transactions carry five to six blobs in a single transaction. In addition, there are some rollups that do not limit how many blobs they carry in a single transaction (such as Linea and Blast).
We also show in 5(b), the percentage of type-3 transactions issued by different rollups that carry the same number of blobs. 69.85 % of the transactions in the network carry only one blob, while a staggering 1,239 transactions carry four blobs. Rollups that send one blob per type-3 transaction dominate. Most rollups use relatively fixed strategies. In addition to that, 2-4 blobs per transaction is a minimal amount.
(a) The blob packing strategies among rollups.
[…]
(b) The landscape of blob usage strategies.
[…]
Block size.
Figure 5. On left is the blob number distribution bar chart of type-3 transactions, on right is the relationship between delays and tips among different rollups.
5.2. Inefficient Strategies
(a) The behaviors and amount of inefficient strategies.
[…]
Block size.
(b) The cumulative loss of inefficient strategies.
[…]
Type-3 excluded block size.
Figure 6. On left is the overview of inefficient strategies from rollups. On right is the cumulative loss of Continuous Sending for rollups.
To analyze whether rollups are using type-3 transactions and blobs efficiently, we analyze their blob-carrying strategies and identify two existing inefficient strategies, Continuous Sending and Double Sending, and another usage, blobscription. 6(a) presents a diagram of these inefficient strategies. A Continuous Sending pattern is when the type-3 transaction sender sends more than one transaction in succession that are ultimately contained within the same block. These blobs can be put into a single type-3 transaction instead of multiple ones, which would require duplicate payments. Double Sending occurs when a rollup transmits the identical blob multiple times. A type-3 transaction that carries a second identical blob is invalid because the first successful transaction completes the validation of the blob data. These two strategies account for 7% and 10% of the total transactions, respectively. Additionally, it is important to mention that blobscription is observed in 106,051 transactions in the network, representing 8% of the total. This section evaluates the impact and economic loss to the network of the inefficient strategies, showing the loss of 186.92 ETH and an average delay in network inclusion of 19 seconds.
Economic Cost.  We calculate the transaction fee (including base fee and priority fee) for duplicate transactions in Continuous Sending and Double Sending.  6(b) plots the cumulative losses from the Double Sending and Continuous Sending strategies for rollups after EIP-4844. The left y-axis shows the cumulative loss for these two strategies, and the right y-axis shows the total number of type-3 transactions. We use light blue bars to show the most rollups per month using these two strategies, and the yellow line represents cumulative losses. Since EIP-4844, rollups have lost 186.92 ETH. while Scroll sends the largest volume of transactions using the two inefficient strategies, it is worth noting that the number of these strategies has been decreasing since May, with less than 1 transaction in August. Another finding is that the use of the continuous sending strategy almost stopped in April, this is because Linea sent significantly fewer type-3 transactions in April.
Network Delay.  The presence of a large number of blobscription transactions violates the original intent of the EIP-4844 design and interferes with the normal rollup validation process. Since a block can only carry a maximum of 6 blobs, the massive number of blobscription transactions in the public mempool may cause a delay in the normal type-3 transactions. We further measure the impact of blobscription on network latency. We analyze 100,504 blobscription transaction during a period of burst of blobscription usage. During this period, there were 151,688 type-3 transaction sent from the rollups, and 21,282 of them were affected by blobscription. The results show that the average latency of the affected type-3 transaction was 43 seconds, while the other type-3 transactions only had to wait 24 seconds. The presence of blobscription in the network introduces a delay of nearly 19 seconds to the rollups. Our finding exposes that if a large number of blobs that are not used for data validation are consistently present within the network, this causes a huge impact on the process of uploading type-3 transactions for rollups, which in turn affects the efficiency of the rollup application.
Answer to RQ2 69.85% of type-3 transactions carried only one blob, with rollups like Taiko and Scroll optimizing for latency, while others like Base and Op Mainnet frequently packed 5–6 blobs per transaction for higher data throughput. Our analysis identifies two inefficient blob-carrying strategies — Continuous Sending and Double Sending—used by rollups, both leading to redundant data submissions and unnecessary costs. Combined with 106,051 blobscription transactions, these inefficiencies caused a total loss of 186.92 ETH and an average network inclusion delay of 19 seconds.
6. Understanding Optimal Strategies
The previous sections have separately examined the trading strategies of builders and rollups in the post-EIP-4844 period. The results indicate that block construction strategies adopted by builders become inefficient with the emergence of type-3 transactions, and that the submission strategies of type-3 transactions through rollups also exhibit inefficiencies. To address this, the interaction between builders and rollups is modeled in this section to determine the optimal pricing of type-3 transactions. The resulting analysis provides both the optimal price that builders should select and the preferred blob-carrying strategy that rollups should adopt.
6.1. Optimal Price for Builders
Table 2. Frequently used notations.
Notation Description
$k$ the number of type-3 transactions in one slot
$q$ the number of blobs carried in one type-3 transaction
$\mathbb{N}$ a set of type-3 transaction $\mathbb{N}=\{n_{1},n_{2},\ldots,n_{k}\}$
$P(n)$ the floor price for blockspace usage by type-3 transactions $n$
$R,\tilde{R}$ the revenue of type-3 transaction and its initial revenue without discount.
$\rho(t)$ the time-dependent discount factor of type-3 transaction revenue.
$b(R)$ the pricing strategy of the type-3 transactions with revenue $R$.
$\pi(R)$ the winning probability of the type-3 transaction with revenue $R$
In the game between builders and rollups, builders privately select a subset of transactions to fill the remaining blockspace, and each rollup submits its type-3 transactions without observing the price of others—mirroring the sealed-bid setting. Moreover, the payment rule is effectively first-price: when a transaction is selected, it pays the exact fee it has given. To capture these mechanics, as well as the effect of limited blockspace and the presence of a floor price $P(n)$, which we have shown in Section 4.2, we formalize the selection process as a first-price sealed-bid auction with size constraints and a reserve floor price $P(n)$. For clarity and ease of reference, Table 2 provides a summary of the notations used throughout the paper, including those that have already appeared in earlier sections.
We assume that type-3 transactions $\mathbb{N}$ compete for the remaining space $s^{*}$ in each block. At this stage, type-3 transactions $\mathbb{N}$ are assumed to compete for the remaining blockspace $s^{*}$. As defined in Section 4.2, a type-3 transaction is represented as $n$. The revenue of type-3 transaction is denoted as $R$. It should be noted that type-3 transactions are time sensitive. For instance, fraud proofs in optimistic rollups must be submitted within a strict challenge window, while validity proofs in ZK-rollups directly determine the finalization of batched Layer-2 transactions on the mainnet (Scroll, 2025; Arbitrum, 2025; Taiko, 2025). In practice, rollups often prioritize certain type-3 transactions—such as validity proofs, fraud proofs, and withdrawal finalization proofs—due to their direct impact on security and user experience. Therefore, a time-dependent discount factor $\rho(t)$ is introduced, where $t$ denotes discrete block slots, i.e., each increment of $t$ corresponds to one block proposal opportunity. The discount captures the fact that the economic value of a type-3 transaction decays as it waits to be included in later slots. Under this formulation, the revenue can be expressed as
(3) $R(q,t)=\rho(t)\tilde{R},128q\leq s^{*},$
where $\tilde{R}$ denotes the revenue at the initial time period $t=0$. The function $\rho(t)$ is strictly decreasing, with $\rho(0)=1$ and $\rho^{\prime}(t)<0$.
Under this setting, each type-3 transaction $n$ must strategically determine its price pp—balancing the inclusion probability against utility maximization. If all type-3 transactions adopt a strictly increasing and differentiable pricing strategy, $p=b(R)$, a type-3 transaction with revenue $R$ wins if and only if its price exceeds all other price.
Further, unlike ordinary auction games, the selection probability of type-3 transactions may also depend on transaction size. Hence, before defining the winning probability, it is necessary to evaluate whether block size constraints affect the inclusion probability of type-3 transactions. Figure 7 presents the relationship between the number of blobs (transaction size) and the average inclusion latency of type-3 transactions after EIP-4844, with a fitted trend for the observed delays. The y-axis represents the inclusion delay, and the x-axis indicates the number of blobs per transaction. Bar colors, ranging from blue to green, correspond to transaction counts. Empirical evidence shows that, after June, type-3 transactions containing between 1 and 5 blobs incur negligible additional latency. Under these conditions, it is reasonable to assume that sufficient blockspace is available, and block size constraints do not materially affect the selection probability. Consequently, the probability of inclusion can be simply expressed as $\pi(R)=F(R)^{n-1}$. The analysis first focuses on a single time period setting, assuming that the revenue $R_{t_{0}}=\rho(t_{0})\tilde{R}$ follows a distribution characterized by the cumulative density function $F(\cdot)$. With the reserve price $P(n)$, the type-3 transaction’s expected utility is
(4) $\mathbb{E}\left[U_{t_{0}}\right]=\Pr\left[\mathrm{win}|R\right]\left(R_{t_{0}}-p\right)=\begin{cases}\pi(R_{t_{0}})\left(R_{t_{0}}-b(R_{t_{0}})\right),\quad&b(R_{t_{0}})\geq P(n).\\ 0,\quad&b(R_{t_{0}})<P(n).\end{cases}$
At symmetric equilibrium, each type-3 transaction selects $p=b(R)$ to maximize $U$. Differentiation with respect to $p$, together with the envelope condition, yields the standard first-order condition (Krishna, 2009) as
(5) $p=\rho(t_{0})\left(\tilde{R}-\frac{\int_{0}^{R_{t_{0}}}\pi(x)dx}{\pi(R_{t_{0}})}\right).$
For tractability, $R$ is always assumed to be independently drawn from a uniform distribution in auction games, i.e., $R\sim U(0,R_{max})$ (Krishna, 2009; Qin et al., 2022). We assume that $k$ and $U(0,R_{max})$ are prior knowledge of builders and rollups. Under this simplifying assumption, the price for type-3 transactions accepted by builders in period $t_{0}$ satisfies
(6) $p_{t_{0}}^{*}=\max\left(P(n_{j}),\frac{k-1}{k}\max_{j}\rho(t_{0})\tilde{R}\right),1\leq j\leq k.$
This represents the builder’s optimal selection strategy when the price of type-3 transactions reach the Bayesian Nash equilibrium. It can be observed that, as the number of type-3 transactions increases, the price obtained by the builders increases and gradually approaches the revenue that the rollups derive from these transactions.
6.2. Blob-carrying Strategy for Rollups
While analyzing the optimal price selection strategy of the builder in Section 6.1, the corresponding pricing behavior of rollups has been implicitly addressed. The discussion now shifts to another challenge in submitting type-3 transactions through rollups: the selection of an appropriate blob-carrying strategy. When multiple blobs need to be submitted within a single period, as described in Section 5.1, the rollup must determine how many blobs are bundled into each type-3 transaction. This corresponds to a classical integer partition problem. In practice, rollups such as Taiko, Scroll, and Starknet tend to split these blobs into several type-3 transactions, whereas Arbitrum and Base choose to aggregate all blobs into a single type-3 transaction. A comparison of these strategies is therefore required. We extend the single-period first-price sealed-bid auction model introduced earlier to a multi-round framework.
[…]
Inclusion delay and blob number.
Figure 7. Inclusion delay and blob number.
The analysis starts with a single rollup submitting two blobs, serving as the foundational scenario for further cases. These blobs generate revenues $R_{1}$ and $R_{2}$. We continue our analysis from the previous single time-period game. After failure, transactions enter the next phase and the blob revenue decays to $\rho(k)$, where $k$ is the number of failed rounds. Therefore, the expected utility of multiple periods becomes a geometric distribution form as
(7) $\mathbb{E}[U]=\frac{\pi(\tilde{R)}}{n}\Phi\left(\pi(\tilde{R})\right),\quad\Phi\left(\pi(\tilde{R})\right)=\sum_{k=1}^{\infty}(1-\pi(\tilde{R}))^{k}\rho(k).$
For convenience, we define
(8) $g(R)=\frac{\pi(R)}{n}R\Phi\bigl{(}\pi(R)\bigr{)}.$
With this notation, the expected utility of bundling two blobs into a single type-3 transaction is $\mathbb{E}[U_{\mathrm{bundle}}]=g(R_{1}+R_{2})$, whereas the expected utility of submitting them separately becomes $\mathbb{E}[U_{\mathrm{split}}]=g(R_{1})+g(R_{2})$. The comparison between these strategies thus reduces to examining the curvature of $g(R)$.
From the second derivative of $g(R)$, it is observed that when $\rho(k)$ decreases slowly, $\pi(R)$ grows rapidly, and $g(R)$ exhibits convexity with $\pi^{\prime\prime}(R)$ remaining positive under a uniform distribution. It means that bundling multiple blobs into a single type-3 transaction is generally the more advantageous strategy. Only in cases where $\rho(k)$ decays sharply—indicating high delay sensitivity and reduced curvature of $\pi(R)$$g(R)$ becomes nearly linear or slightly concave, making a split-submission strategy more beneficial. However, empirical evidence in Figure 7 shows that type-3 transactions carrying multiple blobs do not incur significant additional latency. Hence, bundling multiple blobs into a single type-3 transaction is expected to yield higher utility. Starting from the two-blob case analyzed above, the model extends naturally to the general case of multiple blobs, and the conclusion remains valid. From our landscape analysis in Section 5.1, this is exactly the strategy used by Arbitrarum and Base. In conclusion. from our model and derivation, we give the optimal pricing selection strategy of builders and the optimal blob-carrying strategy of rollups, therefore successfully answer the RQ3.
[…]
7. Related Work
Ethereum Upgrades.  Prior works have explored several upgrades related to Ethereum, including the appearance of ERC-20 tokens (Li et al., 2023; Hu et al., 2024b; Gao et al., 2020; Wu et al., 2023; Xia et al., 2021), emergence of NFT (Huang et al., 2024; Ito et al., 2024; Huang et al., 2023), the London hard fork (Liu et al., 2022; Leonardos et al., 2023), and the transition from PoW to PoS (Heimbach et al., 2023b; Silva et al., 2020; Chi et al., 2024; McLaughlin et al., 2023). After these updates, detection frameworks have been proposed to address the security issues posed by cryptocurrency tokens (Hu et al., 2024a; Victor and Weintraud, 2021; Li et al., 2022). Hu et al. (Hu et al., 2023) developed a model to detect Ethereum fraud, while Zhou et al. (Zhou et al., 2024) examined the use of NFT for arbitrage and airdrop acquisition and revealed that it is driven by a small number of dominant players. Regarding the Ethereum merge event, Heimbach et al. (Heimbach et al., 2023b) quantified and analyzed hard-fork arbitrage in lending protocols. The tokens transferred by arbitrageurs to centralized exchanges were valued at more than \$13 million, and were actually withdrawn from the platform lenders during the Merge of Ethereum. This study complements the analysis of EIP-4844 and provides new observations on network efficiency in the interaction between builders and rollups.
Builder Behaviors.  Several previous works, such as (Öz et al., 2023; Ovezik et al., 2024; Gupta et al., 2023; Wang et al., 2025), investigated builder behaviors in PBS. The landscape exhibited that builders had a significant trend of centralization (Heimbach et al., 2023a). Likewise, Wahrstatter et al. (Wahrstätter et al., 2023) elucidated how vertical integration could potentially extract more value and further centralize block construction. Öz et al. (Öz et al., 2024) illustrated the relationship between the profit and market share, and how private order flow intensified the centralization of the builder market. In addition, Yang et al. (Yang et al., 2024) focused on competition and efficiency in MEV-Boost auctions of builders, and Heimbach et al. (Heimbach et al., 2024) quantified the non-atomic arbitrage utilized by builders to profit in the block-building process. Aside from this, blockchain censorship was defined to measure the censorship of builders in (Wahrstätter et al., 2024; Wang et al., 2023), demonstrating neutrality and security pitfalls. However, no prior works explored builder strategies from the perspective of block sizes and building choices. To the best of our knowledge, we are the first to measure builders’ strategies towards distinct transactions and evaluate their efficiency after EIP-4844, offering crucial insights for builders to refine their block construction.
Rollup Analysis.  Extensive studies on the performance of the rollup ecosystem have revealed crucial insights into its present dynamics and potential future advances (Palakkal et al., 2024; Chaliasos et al., 2024). The process of MEV extraction within Layer-2 rollups has been analyzed by Torres et al. in (Torres et al., 2024), while Avil et al. (Khalil and Rahman, 2024) investigated the arbitrage of NFT that occurs in rollups. Gogol et al. (Gogol et al., 2024b) examine the effects of inscriptions on rollups. They find that the rise in meme transactions leads to reduced gas fees for rollups. Furthermore, Stefanos et al. (Chaliasos et al., 2024) provided insight into zero-knowledge rollups and developed a structured approach for their evaluation. The research presented in (Crapis et al., 2024) introduced an economic framework to examine EIP-4844 and its impact on rollups for reducing the cost of mainnet data. Our research provide the first rollup behavior evaluation and offers our insights on efficient strategies, which is useful for rollups to optimize their strategies in the post-EIP-4844 era.
8. Conclusion
The introduction of EIP-4844 has fundamentally reshaped the economic landscape of Ethereum, creating a new marketplace for data availability. This paper offers the first comprehensive empirical and theoretical analysis of this post-upgrade environment, focusing on the evolving interactions between block builders and rollup operators. Leveraging a novel large-scale dataset encompassing over 319 million transactions, we identified key inefficiencies in current market behavior: builders frequently waste their block space for selecting less profitable type-3 transactions, and rollups often engage in redundant or suboptimal blob submission strategies. These practices result not only in significant direct financial losses, but also in broader network externalities, such as increased inclusion delays and resource congestion.
To address these challenges, we developed a formal game-theoretic framework capturing the strategic tensions at play. From this model, we derived provably efficient strategies: a new pricing rule for builders to ensure profitable block construction, and a bundling principle for rollups that minimizes redundant fees. Our findings contribute to a deeper understanding of the economic implications of Ethereum’s evolving infrastructure and offer actionable guidance to optimize outcomes for both builders and rollup operators.